OpenStudy (anonymous):

Solve 0.9(3x+8)≥1.9-(x+4) Thanks

OpenStudy (anonymous):

First distribute:\[2.7x+7.2 \ge 1.9-x-4\]Combine similar terms and isolate x to one side and a constant to the other:\[2.7x+7.2 \ge -x-2.1\]\[3.7x+7.2 \ge -2.1\]\[3.7x \ge -9.3\]\[x \ge -2.5135\]

OpenStudy (anonymous):

The answer choices are in fraction so I suppose the answer will be \[x \ge-93/37 \] also getting rid of decimals